# Unsupervised and Source-Free Ranking of Biomedical Segmentation Models

### v2

@@ 12.3 Covid_IF External Models Perturbation Sweep
12.4 Cells vs Nuclei Prediction
A key assumption of the proposed consistency based approach is that the source and target tasks align, i.e, if a model was trained for nuclei segmentation then it should only be evaluated for nuclei segmentation. For generalist models like [3] and Cellpose-SAM [51], this assumption may not fully hold. The models are finetuned for both cell and nuclei and have no automated mechanism to control which structure is segmented. In datasets like Covid_IF, a channel can implicitly contain information about both classes. We observed that Cellpose-SAM and occasionally seemingly switch to segmenting nuclei instead of cells (Figs. 5 and 6) particularly in ‘challenging’ sections of the image. It is hard to quantify is this always due to a ‘mode’ switch to nuclei segmentation or just poor cell segmentation performance, however it seems clear that at least in some cases the models segment small objects (much smaller than cells) that clearly correspond to the nuclei information present in the serum channel. This ‘mode’ switch reduces cell segmentation performance, but not consistency scores, as predictions remain internally stable.

### v3

@@ 0.F.3 Covid_IF External Models Perturbation Sweep
0.F.4 Generalist Models: Cells vs Nuclei Prediction
A key assumption of the proposed consistency based approach is that the source and target tasks align, i.e, if a model was trained for nuclei segmentation then it should only be evaluated for nuclei segmentation. For generalist models like [archit_segment_2025] and Cellpose-SAM [pachitariu_cellpose-sam_2025], this assumption may not fully hold in all cases. The models are finetuned for both cell and nuclei and have no automated mechanism to control which structure is segmented. In datasets like Covid_IF, a channel can implicitly contain information about both classes. We observed that Cellpose-SAM and occasionally switch to segmenting nuclei instead of cells (Figs.˜6(a) and 6(b)) particularly in ‘challenging’ sections of the image. It is hard to quantify is this always due to a ‘mode’ switch to nuclei segmentation or just poor cell segmentation performance, however it seems clear that at least in some cases the models segment small objects (much smaller than cells) that clearly correspond to the nuclei information present in the serum channel. This ‘mode’ switch reduces cell segmentation performance, but not consistency scores, as predictions remain internally stable. However, despite some examples of ‘off-task’ segmentation in general Cellpose-SAM and SAM are strong performing cell segmentation models and in the majority of cases are able to stay on task given a good raw input signal. Hence, we can see in Figs.˜4 and 22 that our CMR based ranking of these large generalist models and specialist BMZ model [ouyang_bioimage_2022] strongly correlates with the true target performance ranking.
An important note is that the ranking in Fig.˜4 does not show the general superiority of any one model, but is a target dataset specific ranking of automatic segmentation performance. One factor to keep in mind is that the Covid_IF dataset that we tested ranking on was explicitly not included in the training data of , while inclusion for Cellpose-SAM is uncertain. However, this represents realistic ranking scenarios where models will vary their performance across different target datasets due to differences in source-to-target distribution and model architecture. Hence the need for target specific model ranking as provided by our method.